Two small pitchers and one large pitcher can hold 8 cups of water. one large pitcher minus one small pitcher constitutes 2 cups of a water. How many cups of water can each pitcher hold?
Answer by Earlsdon(6294)

Let the capacity of the large pitcher be L cups and the capacity of the small pitcher be S cups.
1) L+2S = 8 "Two small pitchers (2S) and one large pitcher (L) can hold 8 cups of water."
2) L-S = 2 "One large pitcher (L) minus one small pitcher (S) constitute 2 cups of water."
Solve equation 2) for L---> (L = S+2) and substitue into equation 1) to get:
1a) (S+2)+2S = 8 Solve for S.
1b) 3S+2 = 8 Subtract 2 from both sides.
1c) 3S = 6 Divide both sides by 3.
1d) S = 2 and
L = S+2
L = 2+2 = 4
The large pitcher can hold 4 cups.
The small pitcher can hold 2 cups.